OpenStudy (anonymous):

@satellite73 A recipe that makes 6 servings calls for 1/3 cup of orange juice and 1/2 cup of lemonade. Arlene modifies the recipe so that it can serve 9. How many cups of orange juice does Arlene need? A. 1/6 B.1/3 C. 1/2 D.3

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

help me please

hero (hero):

\[\frac{6 \space servings}{9 \space servings} = \frac{1/3 \space cup }{x \space cup}\]

hero (hero):

\[\frac{6}{9} = \frac{1/3}{x}\] Solve for x

OpenStudy (anonymous):

solve this @Hero

OpenStudy (anonymous):

1/2

hero (hero):

Correct

OpenStudy (anonymous):

thanks @Hero